# ConsRec: Denoising Sequential Recommendation through User-Consistent Preference Modeling

### v1

@@ Abstract
User-item interaction histories are pivotal for sequential recommendation systems but often include noise, such as unintended clicks or actions that fail to reflect genuine user preferences. To address this issue, we propose the User-Consistent Preference-based Sequential Recommendation System (ConsRec), designed to capture stable user preferences and filter noisy items from interaction histories. Specifically, ConsRec constructs a user-interacted item graph, learns item similarities from their text representations, and then extracts the maximum connected subgraph from the user-interacted item graph for denoising items. Experimental results on the Yelp and Amazon Product datasets illustrate that ConsRec achieves a 13% improvement over baseline recommendation models, showing its effectiveness in denoising user-interacted items. Further analysis reveals that the denoised interaction histories form semantically tighter clusters of user-preferred items, leading to higher relevance scores for ground-truth targets and more accurate recommendations. All codes are available at https://github.com/NEUIR/ConsRec.
ConsRec: Denoising Sequential Recommendation through User-Consistent Preference Modeling
Haidong Xin1thanks:   indicates equal contribution., Qiushi Xiong [^1] footnotemark: 1, Zhenghao Liu1thanks:   indicates corresponding author., Sen Mei1, Yukun Yan2, Shi Yu2, Shuo Wang2, Yu Gu1, Ge Yu1, Chenyan Xiong3 1School of Computer Science and Engineering, Northeastern University, China 2Department of Computer Science and Technology, Institute for AI, Tsinghua University, China Beijing National Research Center for Information Science and Technology, China 3Language Technologies Institute, Carnegie Mellon University, United States
1 Introduction
@@
Sequential recommendation systems play a pivotal role in modern digital platforms by dynamically aligning recommendations with evolving user preferences, thereby satisfying user needs and mitigating information overload Wang et al. (2019); Fang et al. (2020). While existing approaches predominantly focus on modeling user behavior through interaction histories, a critical challenge remains unaddressed: real-world interaction sequences inherently contain noise from incidental clicks and transient interests that diverge from user stable preferences. As shown in Figure 1, a user consistently purchasing Lego products might occasionally click on unrelated items like gaming consoles due to momentary curiosity. Such noisy interactions dilute the signal of genuine preferences, leading to suboptimal recommendation results.
Figure 1: Illustration of Our ConsRec Method. ConsRec constructs a user-interacted item graph to denoise items for user preference modeling.
@@
Current methodologies primarily follow two paradigms: item-ID based modeling Kang and McAuley (2018); Sun et al. (2019) and item-content based modeling Hou et al. (2022); Liu et al. (2023). While these approaches learn representations from interaction sequences, they indiscriminately incorporate all items during training, implicitly assuming uniform relevance across interactions. This oversight amplifies noise propagation, as models struggle to distinguish between core preferences and transient outliers Lin et al. (2023b); Zhang et al. (2024a). Recent attempts to address noise either adjust item weights or rely on implicit contrastive learning Yang et al. (2023b), but these methods lack explicit mechanisms to identify and eliminate noise, limiting their interpretability.
To model the consistent preference of users, we propose the User-Consistent Preference-based Sequential Recommendation System (ConsRec), which models user stable preferences by denoising their interacted items. Specifically, ConsRec constructs a user-interacted item graph by collecting all items clicked by a specific user and connecting item pairs with a higher semantic similarity. Then, ConsRec extracts the maximum connected subgraph from the user-interacted item graph to identify the items that represent the consistent preferences of users. These identified items of the maximum connected subgraph are from the densest semantic cluster, effectively helping to model stable preferences.
Our experimental results on Amazon Product and Yelp datasets demonstrate the effectiveness of ConsRec by achieving a 13% improvement over different baseline models. ConsRec can also broaden its effectiveness to different recommendation architectures, showing its generalization ability. Further analysis shows that the retained item sets not only exhibit higher intra-cluster semantic similarity but also better match with user preferences in predicting the next item, confirming the robustness of ConsRec in broadening its effectiveness to different recommendation scenarios. Our work bridges the gap between semantic understanding and graph-based noise filtering, offering a principled solution to enhance sequential recommendation performance.
2 Related Work
@@
Sequential recommendation systems aim to model user behaviors by leveraging user-item interaction histories. Early studies rely on the Markov Chain assumption He and McAuley (2016) and Matrix Factorization techniques Rendle (2010) for item prediction. To better represent user preferences, recent works have primarily focused on ID-based recommendation modeling, which encodes the randomly initialized ID embeddings of user-interacted items using various neural architectures. These architectures include Recurrent Neural Networks (RNNs) Hidasi et al. (2016); Liu et al. (2016); Donkers et al. (2017); Yang et al. (2017), Convolutional Neural Networks (CNNs) Tang and Wang (2018); Yuan et al. (2019); Yan et al. (2019), Graph Neural Networks (GNNs) Chang et al. (2021); Zhang et al. (2022b); Cai et al. (2023), and self-attention mechanisms Kang and McAuley (2018); Sun et al. (2019); Du et al. (2023); Yang et al. (2023a).
@@
Instead of representing items using IDs, existing methods focus more on item content-based recommendation systems, leveraging Pretrained Language Models (PLMs) or Large Language Models (LLMs) to enhance item representations by learning textual semantics from their side information Ding et al. (2021); Yuan et al. (2023); Liu et al. (2023); Chen et al. (2024). These approaches Liu et al. (2023); Chen et al. (2024) encode the textual representations of items and further train language models using user-item interaction signals. TASTE Liu et al. (2023) verbalizes both items and user interaction histories, utilizing the T5 model Raffel et al. (2020) to align user and item representations. Leveraging the emergent abilities of LLMs Huang and Chang (2023); Lin et al. (2023a), HLLM Chen et al. (2024) introduces a hierarchical architecture for sequential recommendation comprising the item LLM and the user LLM.
User interaction sequences often contain noise, such as accidental clicks on irrelevant items Lin et al. (2023b). Existing methods focus on self-supervised denoising, which identifies noise by learning the dependencies among items from user-item interactions Lin et al. (2023b); Zhang et al. (2024a). One research direction adjusts the weight distribution of items within sequences to mitigate the influence of noise, rather than directly excluding noisy items Zhou et al. (2022); Chen et al. (2022); Yang et al. (2023b). The other direction Zhang et al. (2023, 2022a); Sun et al. (2021); Qin et al. (2021) focuses on directly identifying and removing noisy items within interaction sequences. These methods predominantly focus on item ID-based denoising approaches, which require sufficient user-item interaction signals to learn effective user and item representations Lam et al. (2008); Schein et al. (2002); Pan et al. (2019); Zhang et al. (2021). However, they overlook the significant role of text-based semantic information in identifying consistent preferences within the user-interacted item sequence Liu et al. (2023); Yuan et al. (2023).
Figure 2: The Model Architecture of ConsRec.
3 Methodology
This section introduces the User-Consistent Preference-based Sequential Recommendation System (ConsRec). We begin by presenting the preliminaries of the sequential recommendation model (Sec. 3.1). As illustrated in Figure 2, ConsRec employs a filtering mechanism to denoise user-interacted items by identifying user-consistent preferences using item similarities (Sec. 3.2) and then learning tailored item representations for similarity estimation using instruction tuning (Sec. 3.3).
3.1 Preliminary of Sequential Recommendation
@@
$\mathcal{H}\rightsquigarrow M_{\text{Rec}}\rightsquigarrow v_{t},$ (1)
where the user-item interaction history is represented as =$\mathcal{H}=\{v_{1},v_{2},\dots,v_{t-1}\}$. The recommendation model $M_{\text{Rec}}$ estimates the relevance between the user-item interaction history $\mathcal{H}$ and an item $v$ by calculating the ranking probability $P(v|\mathcal{H})$:
$P(v|\mathcal{H})=\text{Softmax}_{v}(\vec{\mathcal{H}}\cdot\vec{v}),$ (2)
where $\cdot$ denotes the dot product operation. Then we follow Liu et al. (2023) to get the representations $\vec{v}$ and $\vec{\mathcal{H}}$ for the candidate item $v$ and the user-item interaction history $\mathcal{H}$.
Item Representation. To fully use the semantic information of items, we verbalize the item $v$ using its identifier and $k$ associated attributes <Attr> through the following template:
$X(v)=\text{id:}v(\text{id})\dots\text{<Attr>}_{k}:v(\text{<Attr>}_{k}),$ (3)
@@
$\vec{v}=\text{T5}(X(v)).$ (4)
User Representation. To represent the user preference, we verbalize the user-item interaction history $\mathcal{H}$ as a concatenation of the text representations (Eq. 3) of the items in $\mathcal{H}$:
$X(\mathcal{H})=X(v_{t-1});\dots;X(v_{1}),$ (5)
where ; denotes the concatenation operation. Then we encode $\mathcal{H}$ into embeddings $\vec{\mathcal{H}}$ using the T5 model, which shares the same parameters with the encoder model in Eq. 4:
$\vec{\mathcal{H}}=\text{T5}(X(\mathcal{H})).$ (6)
3.2 Denoising User-Interacted Items Using User-Consistent Preference
In user-item interactions, users may occasionally display sudden behavioral changes or lose interest in previously preferred items. Such fluctuations introduce noise into the interaction history, complicating accurate user preference modeling.
To tackle this issue, we incorporate a denoising module ($M_{\text{Filter}}$) to identify and retain items that reflect consistent user preferences. Then we train the sequential recommendation module ($M_{\text{Rec}}$) by constructing datasets using the denoised user-item interaction history $\mathcal{H}_{\text{Filter}}$ and the $t$-th step item $v_{t}$. After training, the recommendation module is used to predict the item $v_{t}$ based on $\mathcal{H}_{\text{Filter}}$:
$\mathcal{H}\rightsquigarrow M_{\text{Filter}}\xrightarrow{\mathcal{H}_{\text{Filter}}}M_{\text{Rec}}\rightsquigarrow v_{t},$ (7)
where $\mathcal{H}_{\text{Filter}}\subseteq\mathcal{H}$ represents the subset of items retained to capture consistent preferences of the user. To implement the noise filtering model $M_{\text{Filter}}$, we construct a user-interacted item graph by using the item similarity and then extract the maximum connected component from the graph to obtain the set of denoise items $\mathcal{H}_{\text{Filter}}$.
User-Interacted Item Graph. We first introduce the user-interacted item graph, which is used to distinguish some outlier items that are different from the user’s general preference.
$G_{\mathcal{U}}=(\mathcal{V},\mathcal{E}),$ (8)
where $\mathcal{V}=\{v_{1},\dots,v_{n}\}$ represents all items that have been interacted by user $\mathcal{U}$. The $\mathcal{E}$ is an edge matrix:
=$\mathcal{E}=\{e_{ij},1\leq i\leq n,1\leq j\leq n\},$ (9)
where the element $e_{ij}$ indicates whether the $i$-th node $v_{i}$ and the $j$-th node $v_{j}$ are connected with an edge. The edge is established according to the similarity score $S_{ij}$:
$e_{ij}=\begin{cases}0&\text{if }S_{ij}<\tau,\\ 1&\text{if }S_{ij}\geq\tau,\end{cases}$ (10)
where $\tau$ is the threshold of the similarity. $e_{ij}=0$ indicates that the $i$-th node $v_{i}$ and the $j$-th node $v_{j}$ are not connected, while $e_{ij}=1$ indicates that the nodes are connected. To estimate the similarity between $v_{i}$ and $v_{j}$, we encode all items in the graph node set $\mathcal{V}=\{v_{1},\dots,v_{n}\}$ into embeddings $\{\vec{v}_{1},\dots,\vec{v}_{n}\}$. Subsequently, we calculate the similarity score $S_{ij}$ using cosine similarity:
$S_{ij}=\cos(\vec{v}_{i},\vec{v}_{j}).$ (11)
Then we will introduce how to get the item embedding for similarity calculation in Sec. 3.3.
Denoising User-Interacted Items Using the Maximum Connected Component. To emphasize stronger connections among items, we aim to identify the connected components of the user-interacted item graph $G_{\mathcal{U}}$. Specifically, we perform a Breadth-First Search (BFS) to find a set of disjoint connected components:
$\mathcal{C}=\{C_{1},\dots,C_{l}\},$ (12)
where $l$ is the number of connected components, and each $C_{i}$ is a subgraph of $G_{\mathcal{U}}$. For each connected component $C_{i}$, we compute its size, defined as the number of nodes $|\mathcal{V}(C_{i})|$. The maximum connected subgraph $C_{\text{max}}$ is then selected as:
$C_{\text{max}}=\arg\max_{C_{i}}|\mathcal{V}(C_{i})|.$ (13)
These user-interacted items that belong to the node set $\mathcal{V}(C_{\text{max}})$ of the maximum connected subgraph $C_{\text{max}}$ are retained to form the denoised interaction sequence $\mathcal{H}_{\text{Filter}}$:
$\mathcal{H}_{\text{Filter}}=\{v_{j}\mid v_{j}\in\mathcal{H}\land v_{j}\in\mathcal{V}(C_{\text{max}})\}.$ (14)
This retained interaction sequence $\mathcal{H}_{\text{Filter}}$ is then passed to the recommendation module $M_{\text{Rec}}$ to predict the item at the $t$-th step. More detailed theoretical explanations are presented in Appendix A.7.
3.3 Universal Item Representation Learning for User-Interacted Graph Construction
To obtain item representations that are effective for constructing edges in the graph $G_{u}$, we adopt instruction tuning Chung et al. (2024); Wang et al. (2023) to train language models such as T5 Raffel et al. (2020) based on their textual inputs. This approach differs fundamentally from existing denoising-based methods Yuan et al. (2021a); Zhang et al. (2022a); Yang et al. (2023b); Zhang et al. (2024a), as it leverages supervised signals from different downstream recommendation tasks.
To learn text-based item representations for various recommendation tasks $\mathcal{T}$, we represent each item $v$ only through its textual semantic information. Unlike prior works relying on ID-based representations (Eq. 3), we verbalize the item only using its descriptive attributes and exclude the item ID, enhancing the generalization ability of the item representation model:
$\tilde{X}(v)^{\prime}=\text{<Attr>}_{1}:v(\text{<Attr>}_{1})\dots\text{<Attr>}_{k}:v(\text{<Attr>}_{k}).$ (15)
To fully use the user-item interaction =${\mathcal{H}}=\{v_{1},v_{2},\dots,v_{t-1}\}$, we follow previous work in recommendation model pretraining Sun et al. (2019) and use the Next Item Prediction (NIP) task to train item representations (Eq. 2). Specifically, we verbalize the user-item interaction history $\mathcal{H}$ as a concatenation of verbalized items =$\mathcal{H}=\{v_{1},v_{2},\dots,v_{t-1}\}$:
$\tilde{X}^{\prime}(\mathcal{H})=\tilde{X}(v_{t-1})^{\prime};\dots;\tilde{X}(v_{1})^{\prime},$ (16)
Then we use the instruction to illustrate the Next Item Prediction task: “This is $\mathcal{T}$ dataset. Here is the visit history list of the user: $\tilde{X}(\mathcal{H})^{\prime}$, recommend next item”, aiming to predict the $t$-th step item $v_{t}$.
4 Experimental Methodology

### v2

# LISRec: Modeling User Preferences with Learned Item Shortcuts for Sequential Recommendation

@@ Abstract.
User-item interaction histories are pivotal for sequential recommendation systems but often include noise, such as unintended clicks or actions that fail to reflect genuine user preferences. To address this, we propose Learned Item Shortcuts for Sequential Recommendation (LISRec), a novel framework that explicitly captures stable preferences by extracting personalized semantic shortcuts from historical interactions. LISRec first learns task-agnostic semantic representations to assess item similarities, then constructs a personalized semantic graph over all user-interacted items. By identifying the maximal semantic connectivity subset within this graph, LISRec selects the most representative items as semantic shortcuts to guide user preference modeling. This focused representation filters out irrelevant actions while preserving the diversity of genuine interests. Experimental results on the Yelp and Amazon Product datasets illustrate that LISRec achieves a 13% improvement over baseline recommendation models, showing its effectiveness in capturing stable user interests. Further analysis indicates that shortcut-based histories better capture user preferences, making more accurate and relevant recommendations. All codes and datasets are available at https://github.com/NEUIR/LISRec.
Sequential Recommendation, Task-agnostic Item Representation, User Preference Modeling, Text-based Recommendation
copyright: acmlicensedjournalyear: 2026doi: xxxxxxx.xxxxxxxconference: Proceedings of the 32nd SIGKDD Conference on Knowledge Discovery and Data Mining; August 9-13, 2026; Jeju, Koreaisbn: 979-x-xxxx-xxxx-x/xxxx/xxccs: Information systems Recommender systems
Figure 1. Illustration of Our LISRec Method. LISRec constructs a personalized semantic interaction graph and extracts shortcuts, representing dominant item subsets that provide denoised user representations for preference modeling.
1. Introduction
@@
Sequential recommendation systems play a pivotal role in modern digital platforms by dynamically aligning recommendations with evolving user preferences, thereby satisfying user needs and mitigating information overload (wang2019sequential; fang2020deep). While existing approaches predominantly focus on modeling user behavior through interaction histories, a critical challenge remains unaddressed: real-world interaction sequences inherently contain noise from incidental clicks that diverge from user stable preferences. As shown in Figure 1, a user consistently purchasing LEGO products might occasionally click on unrelated items like gaming consoles due to momentary curiosity. Such noisy interactions dilute the signal of genuine preferences, leading to suboptimal recommendation results.
@@
Current methodologies primarily follow two paradigms: item-ID based modeling (Wang2018self; Sun2019Bert4rec) and item-content based modeling (Hou2022uni; liu2023text). While these approaches learn representations from interaction sequences, they indiscriminately incorporate all items during training, implicitly assuming uniform relevance across interactions. This oversight amplifies noise propagation, as models struggle to distinguish between core preferences and transient outliers (lin2023self; zhang2024ssdrec). Recent attempts to address noise either adjust item weights (wang2021denoising), or rely on implicit contrastive learning (yang2023debiased), but these methods lack explicit mechanisms to identify and eliminate noise, limiting their interpretability.
To address these challenges, we propose Learned Item Shortcuts for Sequential Recommendation (LISRec), a novel framework that explicitly captures stable user preferences by extracting personalized semantic shortcuts from interaction histories. Specifically, LISRec first learns task-agnostic semantic representations to estimate similarities between items, and then constructs a personalized semantic graph using all items previously clicked by the user, based on pairwise item similarities. Next, LISRec identifies a Maximal Semantic Connectivity subset of the personalized semantic graph that best captures the user’s underlying preferences, serving as semantic shortcuts to guide the user behavior modeling for recommendation. By focusing on these highly representative items, LISRec effectively filters out irrelevant interactions while preserving the diversity of the user’s genuine interests.
Our experimental results on the Amazon Product and Yelp datasets demonstrate that LISRec achieves up to 13% improvement over a range of baseline models. Our framework generalizes well across different recommendation architectures, highlighting its flexibility. Further analysis shows that the extracted shortcuts consist of semantically similar items, and that these subsets are more predictive of user preferences in next-item recommendation. Overall, LISRec bridges the gap between semantic representation and noise filtering, providing a principled and generalizable solution for improving sequential recommendation performance.
2. Related Work
@@
Sequential recommendation systems aim to model user behaviors by leveraging user-item interaction histories. Early studies rely on the Markov Chain assumption (he2016fusing) and Matrix Factorization techniques (Rendle2010Factor) for item prediction. To better represent user preferences, recent works have primarily focused on ID-based modeling, which encodes the randomly initialized ID embeddings of user-interacted items using various neural architectures. These architectures include Recurrent Neural Networks (RNNs) (Hidasi2015session; liu2016context; donkers2017sequential; yang2017neural), Convolutional Neural Networks (CNNs) (Jiaxi2018Personalized; Fajie2018asc; yan2019cosrec), Graph Neural Networks (GNNs) (Chang2021graph; zhang2022dynamic; cai2023lightgcl), and self-attention mechanisms (Wang2018self; Sun2019Bert4rec; du2023frequency; yang2023collaborative).
@@
Instead of representing items using IDs, existing methods focus more on item content-based recommendation systems, leveraging Pretrained Language Models (PLMs) or Large Language Models (LLMs) to enhance item representations by learning textual semantics from their side information (Ding2021zero; yuan2023go; liu2023text; chen2024hllm). These approaches (liu2023text; chen2024hllm) encode the textual representations of items and further train language models using user-item interaction signals. TASTE (liu2023text) verbalizes both items and user interaction histories, utilizing the T5 model (raffel2020exploring) to align user and item representations. Leveraging the emergent abilities of LLMs (huang2023towards; lin2023can), HLLM (chen2024hllm) introduces a hierarchical architecture for sequential recommendation comprising the item LLM and the user LLM.
Figure 2. The Model Architecture of LISRec.
User interaction sequences in recommendation systems often contain noise, such as accidental clicks or short-term behaviors that do not reflect users’ true preferences (lin2023self). To address this, several methods reduce the influence of noisy interactions by monitoring training dynamics (e.g., high prediction loss or unstable gradients), including adaptive sample reweighting (wang2021denoising), label correction via double correction (he2024double), and user-specific denoising strategies (zhang2025personalized). Beyond loss-based methods, self-supervised denoising approaches (lin2023self; zhang2024ssdrec) learn item dependencies to identify noisy signals, while other works adjust item weights based on estimated importance (zhou2022filter; chen2022denoising; yang2023debiased). However, these approaches typically treat denoising as an implicit process without directly removing noisy interactions.
Another line of research (zhang2023denoising; zhang2022hierarchical; sun2021does; qin2021world) focuses on explicitly filtering out noise by identifying and removing low-quality items from interaction sequences. These methods predominantly rely on item ID-based representations, which require extensive user-item interaction histories to be effective (lam2008addressing; schein2002methods; pan2019warm; zhang2021language). Consequently, they often overlook the potential of text-based semantic information for identifying preference-consistent behaviors (liu2023text; yuan2023go).
3. Methodology
This section introduces LISRec, a framework for robustly modeling user preferences by extracting semantic shortcuts in a latent space. As shown in Figure 2, we first describe the overall architecture of LISRec, which identifies preference-consistent shortcuts to construct denoised user representations (Sec. 3.1). We then detail the shortcut extraction process, which identifies the Maximal Semantic Connectivity within each user’s personalized item graph (Sec. 3.2).
3.1. The Framework of LISRec
@@
(1) $\small\mathcal{H}\rightsquigarrow M_{\text{Rec}}\rightsquigarrow v_{t},$
where the user-item interaction history is represented as $\mathcal{H}=\{v_{1},v_{2},\dots,v_{t-1}\}$. The recommendation model $M_{\text{Rec}}$ estimates the relevance between the user-item interaction history $\mathcal{H}$ and an item $v$ by calculating the ranking probability $P(v|\mathcal{H})$:
(2) $\small P(v|\mathcal{H})=\text{Softmax}_{v}(\vec{\mathcal{H}}\cdot\vec{v}),$
where $\cdot$ denotes the dot product operation. Then we follow liu2023text to get the representations $\vec{v}$ and $\vec{\mathcal{H}}$ for the candidate item $v$ and the user-item interaction history $\mathcal{H}$.
3.1.1. Item Representation.
To fully use the semantic information of items, we verbalize the item $v$ using its identifier and $k$ associated attributes <Attr> through the following template:
(3) $\small X(v)=\texttt{id:}v(\texttt{id})\dots\texttt{<Attr>}_{k}:v(\texttt{<Attr>}_{k}),$
where $\texttt{<Attr>}_{k}$ is the name of the $k$-th attribute, while $v(\texttt{id})$ and $v(\texttt{<Attr>}_{i})$ correspond to the text descriptions of the item identifier and the $i$-th attribute, respectively. Then we encode each item $v$ into an embedding $\vec{v}$ using the T5 model (raffel2020exploring):
(4) $\small\vec{v}=\text{T5}(X(v)).$
3.1.2. Modeling User Preferences via Shortcut Extraction
Users’ interaction histories ($\mathcal{H}$) often contain a mixture of long-term interests, short-term preferences, and incidental click behaviors. To model user intent using truly essential items, LISRec tries to model preference-consistent behaviors by building semantic shortcuts in the user-interacted item trajectory to omit some noisy items.
To identify these shortcuts, we introduce a semantic shortcut extractor $M_{\text{Filter}}$, which selectively retrieves a subset of dominant interactions from the full history $\mathcal{H}$. These shortcuts, denoted as $\mathcal{H}_{\text{Shortcut}}$, form a condensed yet highly informative sketch of user preferences in the latent semantic space. This distilled representation is then passed to the recommendation module $M_{\text{Rec}}$ for more effective user behavior modeling. The user representation can be formalized as:
(5) $\small\mathcal{H}\xrightarrow{M_{\text{Filter}}}\mathcal{H}_{\text{Shortcut}},$
where $\mathcal{H}_{\text{Shortcut}}\subseteq\mathcal{H}$ denotes the extracted shortcut subset, containing interactions that are semantically aligned with the user’s underlying intent. The semantic shortcut extractor $M_{\text{Filter}}$ are detailed in Section 3.2.
To construct the user representation, we first verbalize the extracted shortcut set $\mathcal{H}_{\text{Shortcut}}$ by concatenating the textual representations (Eq. 3) of the selected items:
(6) $\small X(\mathcal{H}_{\text{Shortcut}})=\mathop{\|}_{v_{i}\in\mathcal{H}_{\text{Shortcut}}}X(v_{i}),$
where $\|$ denotes the concatenation operation over the subset of items in $\mathcal{H}_{\text{Shortcut}}$, which may be non-contiguous and unordered with respect to the original history. The concatenated text is then encoded using the T5 model to obtain a semantic representation of the user’s preference:
(7) $\small\vec{\mathcal{H}}_{\text{Shortcut}}=\text{T5}(X(\mathcal{H}_{\text{Shortcut}})).$
3.2. Extracting Preference Shortcuts via Maximal Semantic Connectivity
To robustly model stable user preferences from noisy interaction histories $\mathcal{H}$, we introduce a semantic shortcut extractor $M_{\text{Filter}}$, which identifies the most representative subset of interactions for preference modeling. Specifically, LISRec first encodes items into a task-agnostic semantic space to estimate pairwise item similarity, then constructs a personalized semantic graph for each user and extracts the maximally connected component as the preference-consistent shortcut. This process effectively filters out noisy or inconsistent behaviors during user representation construction.
3.2.1. Task-Agnostic Semantic Item Encoding
To obtain semantically meaningful item representations suitable for preference-level reasoning, we train a language model on item textual descriptions using the instruction tuning based item representation method.
For item semantic representation, each item $v$ is represented purely through its descriptive attributes—excluding any item ID—to improve generalization and reduce dependency on sparse or domain-specific identifiers:
(8) $\small\tilde{X}(v)^{\prime}=\texttt{<Attr>}_{1}:v(\texttt{<Attr>}_{1})\ \dots\ \texttt{<Attr>}_{k}:v(\texttt{<Attr>}_{k}).$
Building upon the Next Item Prediction (NIP) objective (Sun2019Bert4rec), we serialize each user history $\mathcal{H}=\{v_{1},\dots,v_{t-1}\}$ into a concatenated text sequence:
(9) $\small\tilde{X}^{\prime}(\mathcal{H})=\tilde{X}(v_{t-1})^{\prime};\dots;\tilde{X}(v_{1})^{\prime}.$
@@ 3.2.1. Task-Agnostic Semantic Item Encoding
(10) $\small S_{ij}=\cos(\vec{v}_{i},\vec{v}_{j}).$
This representation approach is fully text-driven, enabling the item representation model to operate across item vocabularies and different tasks while capturing semantic regularities grounded in language.
3.2.2. Modeling Stable Preferences with Personal Item Graphs
To extract the user’s core preference signals from noisy interaction histories, we construct a personalized semantic interaction graph for each user $\mathcal{U}$, where edges capture semantic coherence between interacted items:
(11) $\small G_{\mathcal{U}}=(\mathcal{V},\mathcal{E}),$
where $\mathcal{V}=\{v_{1},\dots,v_{n}\}$ includes all items interacted by user $\mathcal{U}$ in all previous sessions, and $\mathcal{E}$ denotes the edge set encoding semantic connections. An undirected edge $e_{ij}\in\mathcal{E}$ is created between items $v_{i}$ and $v_{j}$ if their embedding similarity exceeds a threshold $\tau$:
(12) $\small e_{ij}=\begin{cases}0&\text{if }S_{ij}<\tau,\\ 1&\text{if }S_{ij}\geq\tau,\end{cases}$
where $e_{ij}=0$ is the cosine similarity between the embeddings of $v_{i}$ and $v_{j}$ are not connected (Eq. 10). This design naturally filters out semantically dissimilar and outlier items with respect to user preferences.
To identify the user’s dominant preference signal, we extract the largest connected component from the graph $G_{\mathcal{U}}$:
(13) $\small C_{\text{max}}=\arg\max_{C_{i}}|\mathcal{V}(C_{i})|,$
and define the corresponding item subset as the preference shortcut:
(14) $\small\mathcal{H}_{\text{Shortcut}}=\{v_{j}\mid v_{j}\in\mathcal{H}\land v_{j}\in\mathcal{V}(C_{\text{max}})\}.$
This shortcut subset serves as a semantically consistent and compact proxy for the user’s intent. By selectively modeling only the maximally connected region in the semantic graph, we avoid overfitting to noisy interactions while mitigating over, ensuring richer and more robust preference representations for downstream recommendation.
4. Experimental Methodology